OpenStudy (anonymous):

what is the derivative of y=(x^3 + 3)(x^4 + 3)

OpenStudy (science0229):

We can use a derivation formula\[(f(x)g(x))'=f'(x)g(x)+f(x)g'(x)\]

OpenStudy (science0229):

In this problem, f(x)=x^3 + 3 and g(x)=x^4 + 3

OpenStudy (science0229):

So, \[((x^3+3)(x^4+3))'=(x^3+3)'(x^4+3)+(x^3+3)(x^4+3)'=3x^2(x^4+3)+4x^3(x^3+3)=3x^6+9x^2+4x^6+12x^3=10x^6+12x^3+9x^2\]

OpenStudy (science0229):

Sorry about that.

OpenStudy (science0229):

I'll continue off from where it was cut off.

OpenStudy (anonymous):

Personally, I think it's easier to distribute it all out.\[y=(x^3+3)(x^4+3)=x^7+3x^4+ 3x^3 + 9\]Now you can just power rule everything and that's it.\[\frac{dy}{dx}=7x^6+12x^3+9x^2\]

OpenStudy (science0229):

Oh yeah

OpenStudy (science0229):

I was going the hard way...

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

@science0229 That's the annoying part about calculus, there's always some simple way to do it but if you don't find it that way, you're probably screwed, lol.

OpenStudy (anonymous):

You're welcome